Which of the following bonds is most likely to be nonpolar covalent bond? P-H Na-Cl Li-Cl O-H C-H
Answer #1

Pauling electronegativity of P = 2.19

Pauling electronegativity of H = 2.20

Difference in electronegativity between P and H = (2.20 - 2.19) = 0.01

Pauling electronegativity of Na = 0.93

Pauling electronegativity of Cl = 3.16

Difference in electronegativity between Na and Cl = (3.16 - 0.93) = 2.23

Pauling electronegativity of Li = 0.98

Pauling electronegativity of Cl = 3.16

Difference in electronegativity between Na and Cl = (3.16 - 0.98) = 2.18

Pauling electronegativity of O = 3.44

Pauling electronegativity of H = 2.20

Difference in electronegativity between Na and Cl = (3.44 - 2.20) = 1.24

Pauling electronegativity of C = 2.55

Pauling electronegativity of H = 2.20

Difference in electronegativity between Na and Cl = (2.55 - 2.20) = 0.35

For a bond to be nonpolar covalent bond, the electronegativity difference between the participating elements should be 0.

The difference in electronegativity between P and H is the least (0.01). Hence, P-H bond is most likely to be non-polar covalent bond.